OpenStudy (anonymous):

Sandra needs 6 pieces of rope, each measuring 8 inches in length.She is going to cut the 6 pieces from a rope that is 84 inches long. How much rope will remain after Sandra cuts the 6 pieces?

OpenStudy (mendicant_bias):

I think you only posted part of the question. www.youtube.com/watch?v=07So_lJQyqw

OpenStudy (anonymous):

Sandra needs 6 pieces of rope, each measuring 8 inches in length. She is going to cut the 6 pieces from a rope that is 84 inches long. How much rope will remain after sandra cuts the 6 pieces?

OpenStudy (anonymous):

36 in.

OpenStudy (mendicant_bias):

Here comes the pitch.....and, oh! Falling just short. So, here's how you would solve this problem: You have a single piece of rope, measuring 84 inches. Let's call that the total length. You need 6 pieces each measuring 8 inches. We'll call the sum of that the cut length. What you need to do to find out the amount you have left over is to subtract the cut length from the total length. How do you find the cut length? You can multiply, like this, since you know you need six pieces, of eight inches length each. \[l _{total} = 84 i n .\]\[l _{cut} = 6 * 8 i n. = 48 i n.\]\[l _{remain} = 84 - 48 = 32 i n.\] Does this make sense?

OpenStudy (mendicant_bias):

Whoops, *36 in.